OpenStudy (anonymous):

If f(x)= -7x+22 then f(2)= and f(x+1)=

OpenStudy (anonymous):

HI welcome to Openstudy! f(x)--> f(2) x=2 so just substitute 2 on the equation f(x)= -7x+22 \(f(2)=-7(2)+22\) same thing for f(x+1) it will be: \(f(x+1)=-7(x+1)+22\) simplify them and tell me what you get.

OpenStudy (anonymous):

Okay f(2) =8 and f(x+1)= -7x-21 or 3

OpenStudy (anonymous):

And thank you for your assistance

OpenStudy (anonymous):

\(f(2)=−7(2)+22\\=-14+22=8\) this seems right:) next one.. \(f(x+1)= -7(x+1)-22\\=-7x-7-22\\=?\)

OpenStudy (anonymous):

no problem:)

OpenStudy (anonymous):

Ohhh I mistakenly placed one where it suppose to be 7--- -7x-29

OpenStudy (anonymous):

Is that right ?

OpenStudy (anonymous):

yes :) sorry i was already off last night